OpenStudy (moonlitfate):

Write the equation for the differential, written dy, for y = sec (5x).Simplify, where possible.

OpenStudy (anonymous):

dy=(dy/dx)dx

OpenStudy (amistre64):

what do you get for y'?

OpenStudy (moonlitfate):

@amistre64 I got y= 5(sec(5x))(tan(5x))

OpenStudy (amistre64):

y' = 5 sec(5x) tan(5x) seems appropriate to me therefore: dy/dx = 5 sec(5x) tan(5x) dy = 5 sec(5x) tan(5x) dx and since y = sec(5x) we might have it out as dy = 5y tan(5x) ... but im not too sure if thats a simplification or not :)

OpenStudy (amistre64):

forgot a dx on the end :)